Assume the balls in the box are numbered 1 through 8, and that an experiment consists of randomly selecting 3 balls one after another without replacement. What probability should be assigned to the event that at least one ball has an odd number?
Answer by scott8148(6628)

number of ways to select 3 out of 8 ___ 8C3 = 56

number of ways to select 3 even balls ___ 4C3 = 4

number of ways to select at least one odd ball ___ 56 - 4 = 52

probability of at least one odd ball ___ 52 / 56 = 13 / 14 = .93 (approx)